a) You purchase one Microsoft June 74 put contract for a premium of $2.37. What is your maximum possible profit given 100 units per contract?

b) An investor buys a call at a price of $6.20 with an exercise price of $57. At what stock price will the investor break even on the purchase of the call?

c) You establish a straddle on Walmart using September call and put options with a strike price of $94. The call premium is $7.70 and the put premium is $8.45. What will be your profit or loss if Walmart is selling for $99 in September?

d) At what stock prices will you break even on the straddle in part c)?
Answer #1

a) Premium on the contract = 2.37*100 = 237

Maximum profit = 74*100 - 237 = 7163

b) Break even on the purchase of the call = 57 +6.2 = 63.2

c) Loss on the transaction = 99 -94 - (8.45+7.7) = -11.15

d) Upper break even level = 94+ (8.45+7.7) = 110.15

Lower break even level = 94 - (8.45+7.7) = 77.85